# Junyu Cao and Mohsen Bayati

### v1

@@ 3.5 Sample Complexity
Combining the sample complexity of horizontal moves and vertical moves, we conclude the total sample complexity of APB.
Theorem 3.14 (Sample complexity of APB)
@@
$H(\delta,\varepsilon;a)=\left(\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))+\max\{\tau^{\rightarrow}(\delta,\varepsilon),\tau^{\rightarrow}(\delta/2,a)\}\tau^{\uparrow}(\delta/2,\varepsilon).$
Theorem 3.14 analyzes the (human) sample complexity in a one-dimensional space, which is in the order of $\tilde{O}(\varepsilon^{-2\kappa})$. The horizontal move is in the logarithm order, so the dominant part comes from the vertical moves in the neighborhood of $\theta^{*}$. If human have the ability to differentiate between any two parameters $\theta_{1}$ and $\theta_{2}$ with probability strictly higher than $1/2+\delta^{\prime}$ as long as $\theta_{1}\neq\theta_{2}$, then the total complexity is in the logarithm order of both confidence $1/\delta$ and precision $1/\varepsilon$.
In a $d$-dimensional space, if naively refining parameters along each dimension to make the parameter within a $\varepsilon$-ball (in $\ell_{p}$-norm), then the total complexity is $dH(\delta/d,\varepsilon/d^{1/p};a)$, which scales as $\tilde{O}(d^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa})$. However, in a high-dimensional space, $d$ can be very large while the true parameter may lie in a low-dimensional manifold. However, in high-dimensional spaces, $d$ can be significant even though the true parameter may lie within a low-dimensional manifold. In the following section, we explore leveraging a supervised learning oracle to reduce sample complexity.
4 Alignment: Human-AI Interaction
In the last section, we discussed the framework of purely using human feedback in a one-dimensional space. In practice, a noisy labeled dataset can usually be acquired easily. A noisy labeled dataset can be potentially utilized for learning the low-dimensional representation $\varphi(\cdot)$ and thus reducing the complexity of the refining step. In this section, we propose a two-stage framework where the first stage learns the low-dimensional representation using noisy labels and the second stage refines the model using human feedback.
@@ 4.1 Two-stage Framework
Suppose the true parameter lies in a $s$-dimensional space where $s\ll d$. Given confidence $\delta$ and precision $\varepsilon$, we introduce a two-stage framework SFT+HC:
@@
    In Stage 1, the algorithm feeds $N_{1}^{HAI}(\delta,\varepsilon)$ noisy labeled data $(\boldsymbol{X},Y)$ to the supervised learning oracle (e.g., Lasso), where $\boldsymbol{X}\in\mathbb{R}^{N_{1}^{HAI}\times d}$ is the design matrix and $Y$ is the noisy response. At the end, we learn the $s$-dimensional embedding $\varphi(\cdot)$;
@@
    In Stage 2, the algorithm asks human to refine the estimator along $s$ dimensions. This stage requires $N_{2}^{HAI}(\delta,\varepsilon)$ data points.
Algorithm 3 SFT+HC (Supervised Fine-Tuning + Human Comparison)
@@
2:Stage 1: Learn the embedding $\varphi(\cdot)$ through SFT oracle trained on data $(\boldsymbol{X},Y)$ with size $N_{1}^{HAI}(\delta,\varepsilon)$;
3:Stage 2: Execute Algorithm 2 (APB) by querying $N_{2}^{HAI}(\delta,\varepsilon)$ samples.
The framework design is grounded in the insight that certain representation learning techniques, like autoencoders or deep learning architectures, are inherently robust to noisy labels (Li et al. 2021, Taghanaki et al. 2021). Stage 1 is dedicated to uncovering underlying patterns and structures within the data. Given the effectiveness of human expertise in model refinement, Stage 2 endeavors to enhance model alignment through targeted sample querying. In what follows, we specifically discuss sparse linear models (Example 2.1), where Stage 1 applies Lasso [^3] OLS and thresholding may serve the same purpose as Lasso. to select important features.
@@ 4.2 Illustration on Linear Models
To demonstrate the sample complexity reduction for the two-stage framework compared to purely trained by supervised learning and provide theoretical justifications, we focus on sparse linear models:
@@
where $\varphi(\cdot)$ projects $X_{i}$ to all important features; $S$ is the index set of all important features; the parameter $\boldsymbol{\theta}^{*}$ satisfies that $|\theta^{*}_{i}|\geq\underline{\beta}_{\Theta}>0$ for all $i\in S$. In the initial stage, we employ Lasso for the selection of significant features, followed by human comparisons for alignment. To assess the sample complexity of our two-stage framework, we will sequentially quantify the size of samples needed for two stages. Firstly, we establish the following standard assumptions for the analysis of Lasso.
@@
For the random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ input in the first stage, we assume the following three conditions holds:
@@
    Mutual incoherence: $\max_{j\in S^{c}}\ \|(\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S})^{-1}\boldsymbol{X}_{S}^{\top}x_{j}\|_{1}\leq\alpha_{1}.$
@@
    Lower eigenvalue: $\lambda_{\min}\left(V_{S}\right)\geq\lambda_{0}>0,$ where $V_{S}=\frac{1}{n}\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S}$.
@@
Assumption 4.2 is standard for the analysis of Lasso. Consider a random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ with i.i.d. $\mathcal{N}(0,1)$ entries for example, we can easily verify that it satisfies mutual incoherence, lower eigenvalue, and $\ell_{\infty}$-curvature condition. Under Assumption 4.2 and 4.2 and based on Corollary 7.22 in Wainwright (2019), we derive the following theorem that bounds the infinite distance between vector $\hat{\theta}_{S}$ and $\theta^{*}_{S}$.
@@
Under Assumption 4.2 and 4.2, suppose that we solve the Lagrangian Lasso with regularization parameter
@@
for $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$. Then the optimal solution $\hat{\theta}$ is unique with its support contained within $S$, and satisfies the $\ell_{\infty}$-error bound
@@
all with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Moreover, there is no false inclusion: The solution has its support set $\hat{S}$ contained with the true support set $S$.
@@
Not limited to the linear function, Lasso can also be applied to the nonlinear function class (Plan and Vershynin 2016) or nonparametric variable selection (Li et al. 2023). Moreover, there is a rich literature of representation learning that can be applied in the first stage (Bengio et al. 2013).
Theorem 4.1 implies that as long as $B_{n}\leq\underline{\beta}_{\Theta}$, the variable selection is consistent. Since $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$, we only need
@@
Thus if there are $n$ noisy samples in which
$n\geq\max\left\{\frac{8\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{32C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}}\right\},$
@@
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\max\left\{\frac{8\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{32C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}},\frac{2\log(4/\delta)}{\zeta^{2}}\right\}=O(\sigma^{2}\log d/\underline{\beta}_{\Theta}^{2}).$
@@
Then, when having more than $H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ noisy labeled data at hand, we can have the correct feature selection with probability at least $1-\delta$. Moreover, if we want to do one more step of refinement to restrict the uncertainty set to be within $\mathfrak{r}$-distance where $\mathfrak{r}<\underline{\beta}_{\Theta}$, then we need $H_{0}(\delta,\varepsilon;\mathfrak{r})$ noisy labeled data. Note that $\mathfrak{r}$, which represents the feasible range for refinement, is the parameter that we can choose. The trade-off exists. A smaller value of $\mathfrak{r}$ demands more noisy labels in the initial stage. However, limiting the feasible region to a smaller area could conserve samples required for human comparisons. As previously noted, when both options significantly deviate from the true model, distinguishing the superior one becomes challenging. This challenge implies a greater need for samples to ascertain the correct answer. Consequently, in certain scenarios, opting for $\mathfrak{r}\leq\underline{\beta}_{\Theta}$ to confine the refinement region may be more optimal. Specifically, we select $\mathfrak{r}$ to minimize the total sample complexity:
$\mathfrak{r}^{*}=\min_{0<\mathfrak{r}\leq\underline{\beta}_{\Theta}}H_{0}(\delta,\varepsilon;\mathfrak{r})+H(\delta,\varepsilon;\mathfrak{r}),$ (4.1)
@@
where $H(\delta,\varepsilon;\mathfrak{r})$ is defined in Theorem 3.14. It is important to emphasize that while the first part depends on the noise $\sigma$, the second part does not rely on $\sigma$ but rather on the detection accuracy $\phi(\tilde{p}(\theta))$. Hence, the optimal $\mathfrak{r}^{*}$ is contingent on the function of detection accuracy. For $\theta$ considerably distant from the true parameter $\theta^{*}$, $\tilde{p}(\theta)$ could potentially approach $1/2$, leading to a substantial value for $\phi(\tilde{p}(\theta))$.
4.3 Complexity of SFT+HC
@@
The optimal value of $\mathfrak{r}^{*}$ can be obtained numerically by enumerating and comparing the objectives in Equation (4.1). To establish an upper bound on the sample complexity of the two-stage framework to analytically show the benefit, we set $\mathfrak{r}=\underline{\beta}_{\Theta}$. In this case, we have
@@
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2})\quad\text{and}\quad H(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}\left(s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa}\right),$
@@
where $\bar{\phi}(a,D)=\max_{a\leq\|\theta-\theta^{*}\|\leq\underline{\beta}_{\Theta}}\phi(\tilde{p}(\theta))$. Theorem 4.3 is a direct conclusion from Theorem 3.14 and 4.1.
Theorem 4.3 (Computational complexity of SFT+HC)
@@
Under Assumption 4.2 and 4.2, by setting $N_{1}^{HAI}(\delta)=H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ and then running Algorithm 2 with parameter range $[-\underline{\beta}_{\Theta},\underline{\beta}_{\Theta}]$, it holds that $\|\hat{\theta}-\theta^{*}\|_{p}\leq\varepsilon$ with probability at least $1-2\delta$. Moreover, under Assumption 3.4.1, the expected sample complexity is
@@
$N_{1}^{HAI}(\delta,\varepsilon)+N_{2}^{HAI}(\delta,\varepsilon)=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa}).$
@@
Theorem 4.3 demonstrates the sample complexity of the two-stage framework SFT+HC. When the distance is measured in two-norm (i.e., $p=2$), the complexity is in the order of $\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\kappa}/\varepsilon^{2\kappa})$. Instead, if exclusively utilizing Lasso for the supervised learning oracle to refine the estimator based on $n$ noisy labeled data, subject to certain regularity conditions, the estimation error can be bounded as $\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq c\sigma\sqrt{s\log d/n}$ with high probability for some constant $c>0$. Equivalently, to ensure that the two-norm distance of the estimation error is within $\varepsilon$, we need to acquire $\tilde{O}(\sigma^{2}s\log d/\varepsilon^{2})$ data points.
@@
Comparing the sample complexity of SFT+HC characterized by Theorem 4.3 for $p=2$, the condition for SFT+HC requiring less sample complexity than SFT is characterized in Proposition 4.4.
@@
For sufficiently small $\varepsilon$, SFT+HC requires less sample complexity than pure SFT for solving the $(\varepsilon,\delta)$-alignment problem when (ignoring the logarithm term)
@@
We call the ratio $\sigma^{2}/s^{\kappa}$ as label noise-to-comparison-accuracy ratio (LNCA ratio) and condition (4.2) as LNCA condition. This condition involves two important parameters: the observational noise $\sigma$ and the convergence rate of the accuracy $\kappa$. When the comparison accuracy is bounded away from 1/2, we have $\kappa=0$ and the condition reduces to $\sigma\gtrapprox\varepsilon$, which holds naturally true. In the worst case where $\kappa=1$, the condition becomes $\sigma^{2}\gtrapprox s$. When sample observational noise is higher, the benefit of the extra human comparison step becomes more apparent.

### v2

@@ 4.5 Sample Complexity
Combining the sample complexity of horizontal moves and vertical moves, we conclude the total sample complexity of MAPB.
Theorem 4.14 (Sample complexity of MAPB)
@@
$H(\delta,\varepsilon;a)=\left(\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))+\tau^{\rightarrow}(\delta/2,\varepsilon)\tau^{\uparrow}(\delta/2,\varepsilon).$
Theorem 4.14 analyzes the (human) sample complexity in a one-dimensional space, which is in the order of $\tilde{O}(\varepsilon^{-2\kappa})$. The horizontal move is in the logarithm order, so the dominant part comes from the vertical moves in the neighborhood of $\theta^{*}$. If humans have the ability to differentiate between any two parameters $\theta_{1}$ and $\theta_{2}$, with the probability strictly higher than $1/2+\delta^{\prime}$ as long as $\theta_{1}\neq\theta_{2}$, then the total complexity is in the logarithm order of both confidence $1/\delta$ and precision $1/\varepsilon$.
In a $d$-dimensional space, if naively refining parameters along each dimension to make the parameter within an $\varepsilon$-ball (in $\ell_{\varrho}$-norm), then the total complexity is $dH(\delta/d,\varepsilon/d^{1/\varrho};a)$, which scales as $\tilde{O}(d^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa})$. However, in high-dimensional spaces, $d$ can be significant even though the true parameter may lie within a low-dimensional space. In the following section, we explore leveraging a supervised learning oracle to reduce sample complexity.
5 Model Alignment: Human-AI Interaction
In the previous section, we discussed the framework of using human feedback in a one-dimensional space. In practice, a noisy labeled dataset usually can be acquired easily. A noisy labeled dataset potentially can be used for learning the low-dimensional representation $\varphi(\cdot)$ and thus for reducing the complexity of the refining step. In this section, we propose a two-stage framework where the first stage learns the low-dimensional representation using noisy labels and the second stage refines the model using human feedback.
@@ 5.1 Two-stage Framework
Suppose the true parameter lies in an $s$-dimensional space where $s\ll d$. Given confidence $\delta$ and precision $\varepsilon$, we introduce a two-stage framework SL+LHF:
@@
    In Stage 1, the algorithm feeds $N_{1}^{HAI}(\delta,\varepsilon)$ noisy labeled data $(\boldsymbol{X},Y)$ to the supervised learning oracle (e.g., Lasso), where $\boldsymbol{X}\in\mathbb{R}^{N_{1}^{HAI}\times d}$ is the design matrix and $Y$ is the noisy response. At the end, we learn the $s$-dimensional embedding $\varphi(\cdot)$.
@@
    In Stage 2, the algorithm asks a human to refine the estimator along $s$ dimensions. This stage requires $N_{2}^{HAI}(\delta,\varepsilon)$ data points.
Algorithm 3 SL+LHF (Supervised Fine-Tuning+Human Comparison)
@@
2:Stage 1: Learn the embedding $\varphi(\cdot)$ through SL oracle trained on data $(\boldsymbol{X},Y)$ with size $N_{1}^{HAI}(\delta,\varepsilon)$;
3:Stage 2: Execute Algorithm 2 (MAPB) by querying $N_{2}^{HAI}(\delta,\varepsilon)$ samples.
The framework design is grounded in the insight that certain representation learning techniques, like autoencoders or deep learning architectures, are inherently robust to noisy labels (Li et al. 2021, Taghanaki et al. 2021). Stage 1 is dedicated to uncovering underlying patterns and structures within the data. Given the effectiveness of human expertise in model refinement, Stage 2 serves to enhance model alignment through targeted sample querying. In what follows, we specifically discuss sparse linear models (Example 3.1), where Stage 1 applies Lasso [^2] OLS and thresholding may serve the same purpose as Lasso. to select important features.
@@ 5.2 Illustration on Linear Models
To demonstrate the sample complexity reduction for the two-stage framework, compared to a framework trained purely by supervised learning, and to provide theoretical justifications, we focus on sparse linear models:
@@
where $\varphi(\cdot)$ projects $X_{i}$ to all important features; $S$ is the index set of all important features; and the parameter $\boldsymbol{\theta}^{*}$ satisfies that $|\theta^{*}_{i}|\geq\underline{\beta}_{\Theta}>0$ for all $i\in S$. In the initial stage, we use Lasso for the selection of significant features, followed by human comparisons for model alignment. To assess the sample complexity of our two-stage framework, we sequentially quantify the size of samples needed for the two stages. First, we establish the following standard assumptions for the analysis of Lasso.
@@
For the random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ input in the first stage, we assume that the following three conditions hold:
@@
    Mutual incoherence: $\max_{j\in S^{c}}\ \|(\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S})^{-1}\boldsymbol{X}_{S}^{\top}x_{j}\|_{1}\leq\alpha_{1};$
@@
    Lower eigenvalue: $\lambda_{\min}\left(V_{S}\right)\geq c_{\min}>0,$ where $V_{S}=\frac{1}{n}\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S}$; and
@@
Assumption 5.2 is standard for the analysis of Lasso. For example, consider a random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ with i.i.d. $\mathcal{N}(0,1)$ entries; we can easily verify that it satisfies mutual incoherence, lower eigenvalue, and $\ell_{\infty}$-curvature condition. Under Assumptions 5.2 and 5.2 and based on Corollary 7.22 in Wainwright (2019), we derive the following theorem that bounds the infinite distance between vector $\hat{\theta}_{S}$ and $\theta^{*}_{S}$.
@@
Under Assumptions 5.2 and 5.2, suppose that we solve the Lagrangian Lasso with regularization parameter
@@
for $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$. Then, the optimal solution $\hat{\theta}$ is unique, with its support contained within $S$, and it satisfies the $\ell_{\infty}$-error bound
@@
all with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Moreover, there is no false inclusion: The solution has its support set $\hat{S}$ contained within the true support set $S$.
@@
Lasso is not limited to the linear function but can also be applied to the nonlinear function class (Plan and Vershynin 2016) or to a nonparametric variable selection (Li et al. 2023). Moreover, there is a rich stream of literature on representation learning that can be applied in the first stage (Bengio et al. 2013).
Theorem 5.1 implies that as long as $B_{n}\leq\underline{\beta}_{\Theta}$, the variable selection is consistent. Because $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$, we only need
@@
Thus, if there are $n$ noisy samples in which
$n\geq\max\left\{\frac{32\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{128C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}}\right\},$
@@
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\max\left\{\frac{32\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{128C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}},\frac{2\log(4/\delta)}{\zeta^{2}}\right\}=O(\sigma^{2}\log d/\underline{\beta}_{\Theta}^{2}).$
@@
Then, when we have more than $H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ noisy labeled data at hand, we can have the correct feature selection with probability at least $1-\delta$. Moreover, if we want to undertake one more step of refinement to restrict the uncertainty set to be within $\mathfrak{r}$-distance, where $\mathfrak{r}<\underline{\beta}_{\Theta}$, then we need $H_{0}(\delta,\varepsilon;\mathfrak{r})$ noisy labeled data. Note that $\mathfrak{r}$, which represents the feasible range for refinement, is the parameter that we can choose. The trade-off exists: A smaller value of $\mathfrak{r}$ demands more noisy labels in the initial stage. However, limiting the feasible region to a smaller area would conserve the samples that are required for human comparisons. As previously noted, when both options significantly deviate from the true model, distinguishing the superior one becomes challenging. This challenge implies a need for more samples to ascertain the correct answer. Consequently, in certain scenarios, opting for $\mathfrak{r}\leq\underline{\beta}_{\Theta}$ to confine the refinement region may be optimal. Specifically, we select $\mathfrak{r}$ to minimize the total sample complexity:
$\mathfrak{r}^{*}=\min_{0<\mathfrak{r}\leq\underline{\beta}_{\Theta}}H_{0}(\delta,\varepsilon;\mathfrak{r})+sH(\delta/s,\varepsilon/s^{1/\varrho};\mathfrak{r}),$ (5.1)
@@
where $H(\delta,\varepsilon;\mathfrak{r})$ is defined in Theorem 4.14 and we measure the distance using $l_{\varrho}$-norm. Note that although the first part depends on the noise $\sigma$, the second part relies not on $\sigma$ but on the detection accuracy $\phi(\tilde{p}(\theta))$. This point is important to emphasize. Hence, the optimal $\mathfrak{r}^{*}$ is contingent on the function of detection accuracy. For $\theta$ considerably distant from the true parameter $\theta^{*}$, $\tilde{p}(\theta)$ could potentially approach $1/2$, leading to a substantial value for $\phi(\tilde{p}(\theta))$.
5.3 Complexity of SL+LHF
@@
The optimal value of $\mathfrak{r}^{*}$ can be obtained numerically by enumerating and then comparing the objectives in Equation (5.1). To establish an upper bound on the sample complexity for the two-stage framework so that we analytically show its benefit, we set $\mathfrak{r}=\underline{\beta}_{\Theta}$. In this case, we have
@@
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2})\quad\text{and}\quad sH(\delta/s,\varepsilon/s^{1/\varrho};\underline{\beta}_{\Theta})=\tilde{O}\left(s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa}\right),$
@@
where $\bar{\phi}(a,D)=\max_{a\leq\|\theta-\theta^{*}\|\leq\underline{\beta}_{\Theta}}\phi(\tilde{p}(\theta))$. Theorem 5.3 is a direct conclusion from Theorems 4.14 and 5.1.
Theorem 5.3 (Computational complexity of SL+LHF)
@@
Under Assumptions 5.2 and 5.2, set $N_{1}^{HAI}(\delta)=H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ and then run Algorithm 2 with parameter range $[-\underline{\beta}_{\Theta},\underline{\beta}_{\Theta}]$; then, it holds that $\|\hat{\theta}-\theta^{*}\|_{\varrho}\leq\varepsilon$ with probability at least $1-2\delta$. Moreover, under Assumption 4.4.1, the expected sample complexity is
@@
$N_{1}^{HAI}(\delta,\varepsilon)+N_{2}^{HAI}(\delta,\varepsilon)=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa}).$
@@
Theorem 5.3 demonstrates the sample complexity of the two-stage framework SL+LHF. When the distance is measured in two-norm (i.e., $\varrho=2$), the complexity is in the order of $\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\kappa}/\varepsilon^{2\kappa})$. Instead, if Lasso is used exclusively for the supervised learning oracle to refine the estimator based on $n$ noisy labeled data (subject to certain regularity conditions), the estimation error can be bounded as $\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq c^{\prime}\sigma\sqrt{s\log d/n}$ with high probability for some constant $c^{\prime}>0$. Equivalently, we need to acquire $\tilde{O}(\sigma^{2}s\log d/\varepsilon^{2})$ data points to ensure that the two-norm distance of the estimation error is within $\varepsilon$.
@@
To compare the sample complexity of SL+LHF as indicated by Theorem 5.3 for $\varrho=2$, we characterize the condition of SL+LHF that requires less sample complexity than SL in Proposition 5.4.
@@
For sufficiently small $\varepsilon$, SL+LHF requires less sample complexity than pure SL for solving the $(\varepsilon,\delta)$-alignment problem (ignoring the logarithm term) when:
@@
We call the ratio $\sigma^{2}/s^{\kappa}$ the label noise-to-comparison-accuracy ratio (LNCA ratio), and we call Condition (5.2) the LNCA condition. This condition involves two important parameters: the observational noise $\sigma$ and the convergence rate of the accuracy $\kappa$. When the comparison accuracy is bounded away from 1/2, we have $\kappa=0$, and the LNCA condition reduces to $\sigma\gtrapprox\varepsilon$, which naturally holds true. In the worst case scenario, where $\kappa=1$, the LNCA condition becomes $\sigma^{2}\gtrapprox s$. When the observational noise of the sample is higher, the benefit of the human comparison step becomes more apparent.
As previously mentioned, our framework can be extended to scenarios where the tasks differ between the initial feature-learning stage and the final implementation stage, provided the low-dimensional representation remains consistent. This is because the sample complexity of the second refinement stage in our algorithm remains unaffected as long as the representation stays unchanged. Meanwhile, Lasso must relearn the varying parameters. Therefore, the sufficient condition outlined in Proposition 5.4 continues to hold.